Mauro Products distributes a single product, a woven basket whose selling price is $21 per unit and whose variable expense is $18 per unit. The company’s monthly fixed expense is $4,200.

Required:

1. Calculate the company’s break-even point in unit sales.

2. Calculate the company’s break-even point in dollar sales. (Do not round intermediate calculations.)

3. If the company's fixed expenses increase by $600, what would become the new break-even point in unit sales? In dollar sales? (Do not round intermediate calculations.)
Answer #1

Contribution margin=Sales-Variable cost  

=(21-18)=$3 per unit

a.Breakeven=Fixed expenses/Contribution margin

=(4200/3)=1400 units

b.Breakeven=(1400*21)=$29400

c.New fixed expenses=(4200+600)=$4800

Hence new breakeven=(4800/3)=1600 units

=(1600*21)=$33600